OpenStudy (anonymous):

What is the solution to the equation -4 + 6(x - 2) = 15?

OpenStudy (♪chibiterasu):

Basically we are finding x so: -4 + 6(x - 2) = 15 -4 + 6x - 12 = 15

OpenStudy (anonymous):

the answer is a fraction

OpenStudy (♪chibiterasu):

Now, add 4 to both sides: -4 + 4 + 6x - 12 = 15 + 4 6x - 12 = 19 Add 12 now: 6x - 12 + 12 = 19 + 12 6x = 31

OpenStudy (♪chibiterasu):

Then the solution will be a decimal which we will convert into a fraction.

OpenStudy (♪chibiterasu):

6x = 31 6x/6 = 31/6 x = 5.16667

OpenStudy (♪chibiterasu):

Not sure what 5.166666 is as a fraction :|